# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 7.1. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
It is important that we work modulo $v_{1}$-torsion to make the above claim. For example, by computing the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 3}\right)$ one can observe that the group $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 3}\right)$ is nonzero in every coweight modulo 4. However, the only classes in coweight $cw\equiv 1\,(4)$ are $v_{1}$-torsion for degree reasons. Moreover, these $v_{1}$-torsion classes are permament cycles of the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 4}\right)$, so their impact on this inductive process is negligible
7.3. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$
To aid in our presentation of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$, we use the notation from Section 6.1 to construct a family of $\text{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules which we organize by coweight modulo 4. As we will see, each module is concentrated in precisely 3 of these coweights, reflecting our observations from Proposition 7.6.
@@ 7.3. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$
For each $i\geq 0$, let $Z_{i}^{\mathbb{R}}$ be the $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module given by the direct sum of the columns of the corresponding table.
@@
There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{R}}$-comodules and $\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
$\frac{\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)}{v_{1}\textup{-torsion}}\cong Z^{\mathbb{R}}_{i}$
@@
The case of $i=1$ was computed in Corollary 7.4, and the case of $i=2$ is presented in Fig. 21, Fig. 22, and Fig. 23. The general case is obtained by applying the functor $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i-1})$ to the cellular filtration of $B_{0}^{\mathbb{R}}(1)$ and computing the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)^{\otimes i})$:
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i-1})\otimes\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\bar{\xi}_{1}],[\bar{\tau}_{1}]\}\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i}).$
The differentials are determined by the attaching maps in the algebraic cell complex $B_{0}^{\mathbb{R}}(1)$ as before, hence Proposition 4.6 implies that $E_{4}=E_{\infty}$. Hidden extensions may be obtained using complex Betti realization to the classical case, linearity over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, and the base change functor $\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C})$ combined with the isomorphism from Proposition 6.9:
$\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k)^{\otimes i}/\rho)\cong\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k)^{\otimes i}).$
The calculations are very similar for all $i$. We give the full argument in the case of $14k+$ 1 and leave the rest to the reader. As an example, we depict $Z_{5}^{\mathbb{R}}$ in Fig. 24, Fig. 25, and Fig. 26.
Let $i=4k$, and consider the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)^{\otimes 4k+1})$. By induction, the $E_{1}$-page is isomorphic to
@@
Recall from Table 3 the contributions from each Atiyah–Hirzebruch filtration to each coweight of the $E_{\infty}$-page of the spectral sequence. In particular, since $Z_{4k}^{\mathbb{R}}$ is trivial in coweight $cw\equiv 3\,(4)$, the $E_{\infty}$-page consist of summands from only two Atiyah–Hirzebruch filtration pieces in coweights $cw\equiv 0,1,3\,(4)$. We proceed by analyzing the $E_{\infty}$-page of the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 4k+1}\right)$ one coweight at a time.
Coweight $cw\equiv 0\,(4)$
@@
There is also a $d_{1}$-differential between big staircases $\Sigma^{2i,i}\mathcal{S}$ in Atiyah–Hirzebruch filtrations 2 and 3 in coweight $cw\equiv 2\,(4)$. The classes in each staircase in Atiyah–Hirzebruch filtration 3 which do not support a $d_{1}$ differential are those which do not support an $h_{0}$-tower. These classes cannot support any further differentials for degree reasons, and so they survive to the $E_{\infty}$-page.
The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 is given by the $h_{1}$-attaching map in $B_{0}^{\mathbb{R}}(1)$. In particular, since the $h_{0}$-towers from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$ do not support $h_{1}$ multiplication, each tower survives to $E_{3}$. None of these $h_{0}$-towers can be the target of a $d_{3}$ differentials for degree reasons, so they also survive to the $E_{\infty}$-page. The $d_{2}$-differential between the big flags $\mathcal{F}_{4i,i,2i}$ in Atiyah–Hirzebruch filtration 2 and 3 in coweight $cw\equiv 0\,(4)$ is determined by the structure given in Definition 6.4. In particular, the classes in the big flag in Atiyah–Hirzebruch filtration 0 which survive to $E_{3}$ are those classes which are not $h_{1}$-divisible. For $i=4k$, these are precisely the $h_{0}$-towers in stems $s\equiv i+4\,(8)$, the $J$-towers in stems $s\equiv i\,(8)$ with Adams filtration $f\geq 1$, and the $(\rho,h_{0})$-tower in Adams filtration $f=0$. For degree reasons, the $d_{3}$-differential is trivial on Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$.
Thus, in coweight $cw\equiv 0$, the $E_{\infty}$-page is given by the $h_{0}$-towers from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0,(4)$ which do not come from the big flag:
$\bigoplus_{k=0}^{2k}\Sigma^{4j-4,4\lfloor(j+1)/2\rfloor-4}H;$
the $h_{0}$-towers, $J$-towers and $(\rho,h_{0})$-tower remaining from the big flag $\mathcal{F}_{4i,i,2i}$ in Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$; and the remaining classes from the big staircase $\Sigma^{2i,i}\mathcal{S}$ in Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 2\,(4)$. Base change to $\mathbb{C}$ and linearity give hidden extensions between the $h_{0}$-towers and $J$-towers from Atiyah–Hirzebruch filtration 0 and the remaining classes from the big staircase in Atiyah–Hirzebruch filtration 3. To be precise, the classes in $\Sigma^{2i,i}\mathcal{S}[3]$ in smallest stem and Adams filtration which survive to $E_{\infty}$ are represented by
@@
and their $\tau^{4}$-multiples. Let $x_{i-4}[0]$ denote the generator of $H\subseteq\mathcal{F}_{4i,i,2i}$ from Atiyah–Hirzebruch filtration 0 in lowest stem and Adams-filtration, and let $x_{i-5}[0]$ denote the generator of $J\subseteq\mathcal{F}_{4i,i,2i}$ from the same Atiyah–Hizebruch filtration. Base change to $\mathbb{C}$ gives hidden extensions
@@
Then, linearity over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ gives us that
@@
Multiplication by $b$ gives the same extensions throughout the $E_{\infty}$-page. Altogether, this gives the summand
@@
Figure 24. $Z_{5}^{\mathbb{R}}$ in $cw\equiv 0\,(4)$.
Coweight $cw\equiv 1\,(4)$
@@
The coweight $cw\equiv 1\,(4)$ portion of the spectral sequence consists of classes coming from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 1\,(4)$ and from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 0\,(4)$. The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all of the classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 0\,(4)$ which are divisible by $h_{0}$. The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all of the classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 0\,(4)$ which support multiplication by $h_{1}$. By observation, we see that every class remaining from the big flag $\mathcal{F}_{4i,i,2i}$ on the $E_{2}$-page is of this kind. Moreover, the filtration forces $E_{3}=E_{\infty}$ on Atiyah–Hirzebruch filtration 2, so the only classes which survive are the bases of the $h_{0}$-towers
$\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor(j+1)/2\rfloor-4}\mathbb{F}_{2}[\tau^{4}].$
The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all of the classes from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 1\,(4)$ which are divisible by $h_{1}$. This can be recovered by appropriately shifting the analogous $d_{2}$-differential in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Fig. 14). The $d_{3}$-differential between Atiyah–Hirzebruch filtrations 0 and 3 is determined by the formula $d_{3}(\rho[3])=(\tau h_{1})[0]$ established in Lemma 7.1. In particular, each remaining class from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 1$ is the target of a differential. Similar to the $d_{2}$-differential just mentioned, this can be recovered by appropriately shifting the analogous $d_{3}$-differential in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Fig. 16). Thus, there are no contributions to the $E_{\infty}$-page in coweight $cw\equiv 1\,(4)$ from Atiyah–Hirzebruch filtration 0. Notice that this forces the classes in the summand $\bigoplus_{j=0}^{2k}\Sigma^{4j-4,4\lfloor(j+1)/2\rfloor-4}\mathbb{F}_{2}[\tau^{4}]$ to be $v_{1}$-torsion, hence the $E_{\infty}$-page is trivial in coweight $cw\equiv 1\,(4)$.
Coweight $cw\equiv 2\,(4)$
@@
The coweight $cw\equiv 2\,(4)$ portion of the spectral sequence consists of classes coming from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$, from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 1\,(4)$, and from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 0\,(4)$. The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all of the classes from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$ which are divisible by $h_{1}$. Notice that while the behavior on the big staircase $\Sigma^{2i,i}\mathcal{S}$ can be recovered from the analogous differential in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Fig. 15), the summand

### v2

@@ 7.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
It is important that we work modulo $v_{1}$-torsion to make the above claim. For example, by computing the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 3}\right)$ one can observe that the group $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 3}\right)$ is nonzero in every coweight modulo 4. However, the only classes in coweight $cw\equiv 1\,(4)$ are $v_{1}$-torsion for degree reasons. Moreover, these $v_{1}$-torsion classes are permament cycles of the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 4}\right)$, so their impact on this inductive process is negligible.
7.3. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$
To aid in our presentation of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$, we use the notation from Section 6.1 to construct a family of $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules which we organize by coweight modulo 4. As we will see, each module is concentrated in precisely 3 of these coweights, reflecting our observations from Proposition 7.4.
@@ 7.3. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$
For each $i\geq 0$, let $Z_{i}^{\mathbb{R}}$ be the $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module given by the direct sum of the columns of the corresponding table.
@@
There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{R}}$-comodules and $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
$\frac{\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)}{v_{1}\textup{-torsion}}\cong Z^{\mathbb{R}}_{i}$
@@
The case of $i=1$ was computed in Corollary 7.3, and the case of $i=2$ is presented in Figure 7.12, Figure 7.13, and Figure 7.14. The general case is obtained by applying the functor $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i-1}\otimes-)$ to (4.3) and using the resulting $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)^{\otimes i})$:
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i-1})\otimes\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\bar{\xi}_{1}],[\bar{\tau}_{1}]\}\implies\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i}).$
The differentials are determined by the attaching maps in the algebraic cell complex $B_{0}^{\mathbb{R}}(1)$ as before, hence Proposition 4.4 implies that $\textup{E}_{4}=\textup{E}_{\infty}$. Hidden extensions may be obtained using complex Betti realization to the classical case, linearity over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, and the base change functor $\otimes\mathbb{C}:\text{SH}(\mathbb{R})\to\text{SH}(\mathbb{C})$ combined with the isomorphism from Proposition 6.5:
$\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k)^{\otimes i}/\rho)\cong\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k)^{\otimes i}).$
The calculations are very similar for all $i$. We give the full argument in the case of $4k+1$ and leave the rest to the reader. As an example, we depict $Z_{5}^{\mathbb{R}}$ in Figure 7.16, Figure 7.17, and Figure 7.18.
Let $i=4k$, and consider the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)^{\otimes 4k+1})$. By induction, the $\textup{E}_{1}$-page is isomorphic to
@@
Recall from Table 3 the contributions from each Atiyah–Hirzebruch filtration to each coweight of the $\textup{E}_{\infty}$-page of the spectral sequence. In particular, since $Z_{4k}^{\mathbb{R}}$ is trivial in coweight $cw\equiv 3\,(4)$, the $\textup{E}_{\infty}$-page consist of summands from only two Atiyah–Hirzebruch filtration pieces in coweights $cw\equiv 0,1,3\,(4)$. We proceed by analyzing the $\textup{E}_{\infty}$-page of the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 4k+1}\right)$ one coweight at a time.
$\textup{E}_{\infty}$-coweight $cw\equiv 0\,(4)$
@@
We depict this differential in Figure 7.15.
Figure 7.15. The $d_{1}$-differential between the $H$-towers from Atiyah–Hirzebruch filtrations 3 to 2 in coweight $cw\equiv 2\,(4)$.
There is also a $d_{1}$-differential between big staircases $\Sigma^{2i,i}\mathcal{S}$ in Atiyah–Hirzebruch filtrations 2 and 3 in coweight $cw\equiv 2\,(4)$. The classes in each staircase in Atiyah–Hirzebruch filtration 3 which do not support a $d_{1}$-differential are those which do not support an $h_{0}$-tower. These classes cannot support any further differentials for degree reasons, and so they survive to the $\textup{E}_{\infty}$-page (see Figure 7.3).
The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 is given by the $h_{1}$-attaching map in $B_{0}^{\mathbb{R}}(1)$. Since the $h_{0}$-towers from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$ are $h_{1}$-torsion, each tower survives to $\textup{E}_{3}$. None of these $h_{0}$-towers can be the target of a $d_{3}$-differential for degree reasons, so they also survive to the $\textup{E}_{\infty}$-page.
The $d_{2}$-differential between the big flags $\mathcal{F}_{4i,i,2i}$ in Atiyah–Hirzebruch filtrations 0 and 2 in coweight $cw\equiv 0\,(4)$ is determined by the structure given in Definition 6.3. The classes in the big flag in Atiyah–Hirzebruch filtration 0 which survive to $\textup{E}_{3}$ are those classes which are not $h_{1}$-divisible. For $i=4k$, these are precisely the $h_{0}$-towers in stems $s\equiv i+4\,(8)$, the $J$-towers in stems $s\equiv i\,(8)$ with Adams filtration $f\geq 1$, and the $(\rho,h_{0})$-tower in Adams filtration $f=0$. For degree reasons, the $d_{3}$-differential is trivial on Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$ (see Figure 7.4).
Thus, in coweight $cw\equiv 0\,(4)$, the $\textup{E}_{\infty}$-page is given by the $h_{0}$-towers from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$ which do not come from the big flag:
$\bigoplus_{j=1}^{k+1}\Sigma^{4j-4,4\lfloor(j+1)/2\rfloor-4}H;$
the $h_{0}$-towers, $J$-towers and $(\rho,h_{0})$-tower remaining from the big flag $\mathcal{F}_{4i,i,2i}$ in Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$; and the remaining classes from the big staircase $\Sigma^{2i,i}\mathcal{S}$ in Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 2\,(4)$.
Base change to $\mathbb{C}$ and linearity give hidden extensions between the $h_{0}$-towers and $J$-towers from Atiyah–Hirzebruch filtration 0 and the remaining classes from the big staircase in Atiyah–Hirzebruch filtration 3. To be precise, the classes in $\Sigma^{2i,i}\mathcal{S}[3]$ in smallest stem and Adams filtration which survive to $\textup{E}_{\infty}$ are represented by
@@
and their $\tau^{4}$-multiples. Let $x_{i-4}[0]$ denote the generator of $H\subseteq\mathcal{F}_{4i,i,2i}$ from Atiyah–Hirzebruch filtration 0 in lowest stem and Adams-filtration, and let $x_{i-5}[0]$ denote the generator of $J\subseteq\mathcal{F}_{4i,i,2i}$ from the same Atiyah–Hirzebruch filtration. Base change to $\mathbb{C}$ gives hidden extensions
@@
Then, linearity over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ gives us that
@@
Multiplication by $b$ gives the same extensions throughout the $\textup{E}_{\infty}$-page. Altogether, this gives the summand
@@
The reader may benefit from consulting the chart in Figure 7.9.
As an example, Figure 7.16 depicts the coweight $cw\equiv 0\,(2)$ portion of $Z_{5}^{\mathbb{R}}$. In this case, we have:
  • the $h_{0}$-towers in stems 0 and 4 come from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 0\,(4)$;
  • the $J$-tower in stem 8 is the $v_{1}$-torsion free component of the $(\rho,h_{0})$-tower in Atiyah–Hirzebruch filtration 0;
  • the staircase $\mathcal{S}\langle 1\rangle$ in stem 12 consists of the $h_{0}$-towers and $J$-towers remaining from the big flag in Atiyah–Hirzebruch filtration 0 and the remainder of the big staircase from Atiyah–Hirzebruch filtration 3, together with hidden extensions.
Figure 7.16. $Z_{5}^{\mathbb{R}}$ in $cw\equiv 0\,(4)$.
$\textup{E}_{\infty}$-coweight $cw\equiv 1\,(4)$
@@
The coweight $cw\equiv 1\,(4)$ portion of the spectral sequence consists of classes coming from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 1\,(4)$ and from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 0\,(4)$. The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all of the classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 0\,(4)$ which are divisible by $h_{0}$ (see Figure 7.1 for the differential on the big flag summands and Figure 7.15 for the differentials on the $h_{0}$-towers). The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all of the classes from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 0\,(4)$ which support multiplication by $h_{1}$ (see Figure 7.4). By observation, we see that every class remaining from the big flag $\mathcal{F}_{4i,i,2i}$ on the $\textup{E}_{2}$-page supports $h_{1}$-multiplication. Moreover, the Atiyah–Hirzebruch on $B_{0}^{\mathbb{R}}(1)$ filtration forces $\textup{E}_{3}=\textup{E}_{\infty}$ on Atiyah–Hirzebruch filtration 2, so the only classes which survive are the bases of the $h_{0}$-towers
(7.6) $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor(j+1)/2\rfloor-4}\mathbb{F}_{2}[\tau^{4}].$
@@
The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all of the classes from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 1\,(4)$ which are divisible by $h_{1}$. This can be recovered by appropriately shifting the analogous $d_{2}$-differential in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Figure 7.5). The $d_{3}$-differential between Atiyah–Hirzebruch filtrations 0 and 3 is determined by the formula
$d_{3}(\rho[3])=(\tau h_{1})[0]$
established in Lemma 7.1. In particular, each remaining class from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 1$ is the target of a differential. Similar to the $d_{2}$-differential just mentioned, this can be recovered by appropriately shifting the analogous $d_{3}$-differential in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Figure 7.7). Thus, there are no contributions to the $\textup{E}_{\infty}$-page in coweight $cw\equiv 1\,(4)$ from Atiyah–Hirzebruch filtration 0. For degree reasons, the classes in the summand (7.6) are be $v_{1}$-torsion, hence the $\textup{E}_{\infty}$-page is trivial in coweight $cw\equiv 1\,(4)$.
$\textup{E}_{\infty}$-coweight $cw\equiv 2\,(4)$
The coweight $cw\equiv 2\,(4)$ portion of the spectral sequence consists of classes coming from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$, from Atiyah–Hirzebruch filtration 2 in coweight $cw\equiv 1\,(4)$, and from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 0\,(4)$. The $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 kills all of the classes from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$ which are divisible by $h_{1}$. Notice that while the behavior on the big staircase $\Sigma^{2i,i}\mathcal{S}$ can be recovered from the analogous differential in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Figure 7.6), the summand of $H$-towers
(7.7) $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,4\lfloor j/2\rfloor-2}H$
in Atiyah–Hirzebruch filtration 0 is entirely $h_{1}$-torsion, hence survives to $\textup{E}_{3}$. The $d_{3}$-differential between Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$ and Atiyah–Hirzebruch filtration 3 $cw\equiv 1\,(4)$ can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Figure 7.8). Notice that there can be no $d_{3}$ differentials whose target are the $h_{0}$-towers in (7.7) for degree reasons. This determines the contributions to the $\textup{E}_{\infty}$-page in coweight $cw\equiv 2\,(4)$ coming from Atiyah–Hirzebruch filtration 0.
The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 in coweight $cw\equiv 1\,(4)$ can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Figure 7.2). Similarly, the $d_{2}$-differential between Atiyah–Hirzebruch filtrations 0 and 2 in coweight $cw\equiv 1\,(4)$ can be recovered from the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ (see Figure 7.5). For degree reasons, we have that $\textup{E}_{3}=\textup{E}_{\infty}$ on Atiyah–Hirzebruch filtration 2, so this determines the contributions to the $\textup{E}_{\infty}$-page in coweight $cw\equiv 2\,(4)$ from Atiyah–Hirzebruch filtration 2.
The $d_{1}$-differential between Atiyah–Hirzebruch filtrations 2 and 3 kills all classes from Atiyah–Hirzebruch filtration 3 in coweight $cw\equiv 0\,(4)$ which support $h_{0}$-multiplication. In particular, this kills the summand $\bigoplus_{j=1}^{2k}\Sigma^{4j-4,r\lfloor(j+1)/2\rfloor-4}H$ and all $h_{0}$-towers in the big flag $\mathcal{F}_{4i,i,2i}$. The $d_{3}$-differential between Atiyah–Hirzebruch filtrations 0 and 3 is determined by the formula
$d_{3}(\rho[3])=(\tau h_{1})[0]$
established in Lemma 7.1. In particular, using the notation of Definition 6.3, we have (see Figure 7.7)
$d_{3}(\rho\cdot x_{4k}[3])=(\Sigma^{2i,i}\tau h_{1})[0].$
Linearity gives the rest of the differentials. Note that this is the differential which ensures that Atiyah–Hirzebruch filtration 0 is trivial in coweight $cw\equiv 1\,(4)$ on the $\textup{E}_{\infty}$-page.
Thus, in coweight $cw\equiv 2\,(4)$, the $\textup{E}_{\infty}$-page is given by the remaining $h_{0}$-towers from the big staircase $\Sigma^{2i,i}\mathcal{S}$ from Atiyah–Hirzebruch filtration 0 in coweight $cw\equiv 2\,(4)$ and the $h_{0}$-towers